MICROWAVE SPECTRA AND STRUCTURES OF H_{2}S-CuCl AND H_{2}O-CuCl.

\underline{N. R. WALKER}, D. E. WHEATLEY, S. L. STEPHENS, F. J. ROBERTS, V. A. MIKHAILOV AND A. C. LEGON, School of Chemistry, University of Bristol, Bristol, BS8 1TS, U.K..

A Balle-Flygare FT-MW spectrometer coupled to a laser ablation source has been used to measure the pure rotational spectra of H_{2}S-CuCl and H_{2}O-CuCl. Both molecules are generated via laser ablation (532 nm) of a metal rod in the presence of CCl_{4}, argon, a low partial pressure of H_{2}S or H_{2}O and are stabilized by supersonic expansion. Rotational constants and centrifugal distortion constants have been measured for eight isotopologues of H_{2}S-CuCl with substitutions available at the copper, chlorine and hydrogen atoms. Transitions in the spectra of nine isotopologues of H_{2}O-CuCl have been measured with isotopic substitutions achieved for every atom. The spectra of both H_{2}S-CuCl and H_{2}O-CuCl are consistent with a linear arrangement of sulphur or oxygen, metal and chlorine atoms. The structure of H_{2}S-CuCl is pyramidal with C_{S} symmetry. The structure of H_{2}O-CuCl is either C_{2v} planar at equilibrium or C_{S} pyramidal but with a low potential-energy barrier to planarity such that the v=0 and 1 states associated with the motion that inverts the configuration at the O atom are well separated. Nuclear quadrupole coupling constants have been measured for the chlorine and copper atoms in each molecule. Nuclear spin-rotation constants have been determined for the copper atom.